OpenStudy (jamesreed30):

What is the radius of a circle with a sector area of seven pi square feet and in art whose measure is 70°

KjSaif (kjsaif):

7pi=(7/36)pir^2 7pi=(7pi/36)r^2 252pi=7pir^2 36pi=pir^2 36=r^2 6=r

KjSaif (kjsaif):

@Jamesreed30

OpenStudy (jamesreed30):

Thank you! I've got a couple more